mikewwe13:

Nadia swims at a rate of 50 meters per minute. Create a function f, where f(n) gives the number of meters Nadia swims given the number of minutes she swims, n. Select the options to the empty boxes to correctly complete the equation for the function. f(n) = 50 ____ ____ f n + - ● ÷ Some time ago, Carmen put an initial deposit of $150 into an account. After his initial deposit, he added more money to his account every two weeks. The table shows the amount of money in Carmen's account starting with his initial deposit. Part A. How did the amount of money, f(w), in Carmen's account change every two weeks? Select from the drop-down menus to correctly complete the sentence about the money in Carmen's account. After Carmen's initial deposit, the amount of money his account _______________ increased or decreased at a _____________ varying or constant rate. Part B. Create the function f, where f(w) models the amount of money in Carmen's account given the number of weeks, w, after his initial deposit. f(w) = _______ ● w ________ _________ 40 80 150 450 + - ● ÷

Vocaloid:

if minutes she swims = n how would you express "50 times minutes she swims" in ~~~math~~~ NOT words?

mikewwe13:

1(4 +46 ) = 50 To walk 50 meters in per minute she has to walk 80 feet a minute.

Vocaloid:

please do not put unrelated information from other problems

mikewwe13:

i didn't i just worked this out

Vocaloid:

the problem has nothing to do with walking

mikewwe13:

i meant to say swim but i put walk by accident

Vocaloid:

one more time if n = minutes then how can we express "50 times minutes" in math not words?

Vocaloid:

hint: "times" means multiplication

mikewwe13:

50 x 50

Vocaloid:

try one more time

Vocaloid:

n = minutes 50 * minutes = ?

mikewwe13:

7

Vocaloid:

???

Vocaloid:

n = minutes means we can replace "minutes" with "n" when we see minutes so 50 * minutes = ?

mikewwe13:

n

mikewwe13:

50 x minutes = n

Vocaloid:

nope, one more time

Vocaloid:

write down "50 x minutes" erase "minutes" write "n" where minutes used to be

Vocaloid:

let me know what you get.

mikewwe13:

50 x minutes erases minutes and put n where minutes used to 50 x n

Vocaloid:

good so 50 x n is your answer for A

mikewwe13:

oh i see it

Vocaloid:

for the next problem I need to see the table to answer it.

mikewwe13:

but wait there's no x

Vocaloid:

● is the same as x

mikewwe13:

oh ok

Vocaloid:

both of these symbols represent multiplication

mikewwe13:

ok

Vocaloid:

again, please post a screenshot of the table so I can work on the problem.

mikewwe13:

ok hold on

mikewwe13:

1 attachment
mikewwe13:

that's the table

Vocaloid:

ok let's start by calculating the rate of change in f(w)

Vocaloid:

we use a formula called the rate of change formula rate of change = [ f(x2) - f(x1) ] / (x2-x1)

Vocaloid:

now, let's look at the table where w = 0 and w = 2 what are the values of f(0) and f(2)? look at the table please.

mikewwe13:

on the time, w, in weeks

Vocaloid:

when w = 0, what is the value of f(0)? use the table to answer. your answer should be a number, not words.

mikewwe13:

f (0) = 0

Vocaloid:

look at the table, one row at a time. the left column tells you w the right column tells you f(w) for that same w for example, if you look at the last column, w = 8 and f(w) = f(8) = 470 now, use this same logic to tell me what f(0) is

mikewwe13:

w = 0 and f(w) = 150 = f(0) = 150

Vocaloid:

good

Vocaloid:

what about f(2)?

mikewwe13:

w = 2 and f(w) = 230 = f(2) = 230

Vocaloid:

good so if f(2) = 230 and f(0) = 150 calculate: [ f(2) - f(0) ] / (2 - 0) be careful with parentheses/order of operations

mikewwe13:

f(4)

mikewwe13:

am i correct ?

Vocaloid:

your answer should be a number not including f

Vocaloid:

f(2) = 230 and f(0) = 150

Vocaloid:

replace "f(2)" with 230 and "f(0)" with 150 so calculate [ f(2) - f(0) ] / (2 - 0) = ?

mikewwe13:

230 - 150 ? solve ? like that

Vocaloid:

yes, but for the entire expression

Vocaloid:

[ f(2) - f(0) ] / (2 - 0) = (230-150)/2 = ?

mikewwe13:

40

Vocaloid:

good so the weekly rate of change is 40 (to speed things up, I will tell you that the table has a constant rate of change of 80, you can verify this on your own time using the same formula)

Vocaloid:

now, we know our "intercept" or intial value is f(0) = 150 so we can now write an equation like so:

Vocaloid:

f(w) = (weekly rate of change) * w + initial value replace "weekly rate of change" with 40 and "initial value" with 150 re-write the equation now.

Vocaloid:

BRB

mikewwe13:

ok

mikewwe13:

f(w) = 40 x w + 150

Vocaloid:

Good that's your answer for the problem

mikewwe13:

Wait is Carmen's account increased or decreased ?

mikewwe13:

and varying or constant rate ?

Vocaloid:

Is the money going up or down?

mikewwe13:

down

Vocaloid:

Look at the table.

mikewwe13:

i meant up i meant sorry

Vocaloid:

good so the money is increasing

mikewwe13:

and what about the rate ?

Vocaloid:

I already said it was constant

mikewwe13:

oh wait ok

mikewwe13:

could you help me with another one ?

Vocaloid:

have to get going, sorry

mikewwe13:

please right quick

mikewwe13:

Part A There is a function f, where f(n) = 1200n + 3500 represents the number of people in a town who owned a smartphone n years after 2007. How many people owned a smartphone in this town in the year 2010 ? Enter your answer in the box ____________ Part B There is a function p, where p(x) represents the percent of adults in the United States who owned a smartphone x years after 2011. What does p(5) = 68 represent? Select from the drop-down menus to correctly complete the sentence. In the year ___________ 2011 or 2016 ____________ 5 or 68 or 5% or 68% of the adults in the United States owned a smartphone.